OpenStudy (anonymous):

MEDAL GIVEN: The Parent Teacher Organization has raised $1800 to help pay for a year-end trip to an amusement park. They ask that there be one adult for every five students attending. Adult tickets cost $45 and student tickets cost $30. If the group wants to take 50 students, how much will each student need to pay so that adults agreeing to chaperone pay nothing?

OpenStudy (anonymous):

K give me a sec

OpenStudy (anonymous):

They want to take 50 student? Or 50 ppl

OpenStudy (anonymous):

Students*

OpenStudy (anonymous):

50 students means 50 * 30 = 1500 dollars for student tickets 50 / 5 = 10 adults required 10 * 45 = 450 dollars for adult tickets 1500 + 450 = 1950 dollars needed 1950 / 50 = each student must pay 39 dollars!

OpenStudy (anonymous):

Wait, does the $1800 raised mean the students don't have to pay that amount? IF so, 1950 - 1800 = 150 dollars needed 150 / 50 = 3 dollars needed from each student after using the 1800 raised